a company in the wholesale trade selling sportswear and stocks two brands, a and b, of football kit, each consisting of a shirt, a pair of shorts and a pair of socks. the costs for brand a are sh.50 for a shirt, sh 30 for a pair of shorts and sh10 for a pair of socks and those for brand b are sh.60.25 for a shirt, sh.40 for a pair of shorts and sh.10 for a pair of socks. three customers x, y, z demand the following combinations of brands; x, 36 kits of brand a and 48 kits of brand b y, 24 kits of brand a and 72 kits of brand b z, 60 kits of branda required:i) express the costs of brands a and b in matrix form, then the demands of the customers, x,y and z in matrix form. (5 marks)ii) by forming the product of the two matrices that you obtain in the previous part, deduce the detailed costs to each of the customers. (5 marks)


1
Expert's answer
2021-10-28T16:13:01-0400

(i)

The costs of brands a and b:


shirt shorts socks

brand a 50 30 10

brand b 60.25 40 10


The demands of the customers:


"\begin{matrix}
& \textrm{Brand \ A} & \textrm{Brand \ B} \\
\textrm{Customer \ X} & 36 & 48 \\
\textrm{Customer \ Y} & 24 & 72 \\
\textrm{Customer \ Z} & 60 & 0 \\
\end{matrix}"


(ii)

Detailed costs to each of the customers:


"\begin{pmatrix}
 36 & 48 \\
 24 & 72\\
60&0
\end{pmatrix}\begin{pmatrix}
 50 & 30&10 \\
 60.25 & 40&10
\end{pmatrix}=\begin{pmatrix}
 4692 & 3000& 840 \\
 5538 & 3600& 960 \\
3000 & 1800&600
\end{pmatrix}"


"\begin{matrix}
& \textrm{ shirt} & \textrm{shorts}& \textrm{socks} \\
\textrm{Customer \ X} & 4692 & 3000&840 \\
\textrm{Customer \ Y} & 5538 & 3600&960 \\
\textrm{Customer \ Z} & 3000 & 1800&600 \\
\end{matrix}"